OpenStudy (anonymous):

Which of the following is the solution to the equation 25^2c = square root of 5^(4c - 10) ? A. c = -5/2 B. c = 7/2 C. c = 4 D. c = 7

OpenStudy (anonymous):

\[25^(2c)=\sqrt{5}^(4c-10)\]

OpenStudy (anonymous):

52^2c=5(1/2)×4c−10 4c=2c−5 2c=−5 c=−5÷2 This is what I make of the question. I hope it helps.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yep :)